# Retrieve, Merge, Predict: Augmenting Tables with Data Lakes (Experiment, Analysis & Benchmark Paper)

### v3

@@ 3. Building Yet Another Data Lake
We detail the construction of YADL (Yet Another Data Lake), a synthetic data lake built starting from the YAGO (Suchanek et al., 2007; Mahdisoltani et al., 2014) knowledge base (KB). Our goal is a scalable and reliable data lake that allows users to evaluate the main steps of the pipeline in Fig. 1, while avoiding some of the confounding factors that come from working with unvetted data, such as typos, inconsistent schemas and data format, and other sources of noise. These challenges can easily be added to YADL, e.g., by generating typos in table entries.
3.1. YAGO: the source of the original data
@@
YAGO (Suchanek et al., 2007) is a KB composed of RDF triplets; each triplet has a subject connected to an object through a predicate (or relation). Subjects and objects are considered to be entities, e.g., in the triplet “Paris, locatedIn, France”, “Paris” is the subject entity, “France” is the object entity, and “locatedIn” is the predicate that connects them. The object may also be a lexical value such as the “population density”. Each entity belongs to a set of classes (or types), arranged in a taxonomy; “Paris” belongs to the class “City”, subclass of “Populated place”, itself subclass of “Geographical locations”. We use YAGO 3.0.3 (Mahdisoltani et al., 2014), which is updated to 2022.
3.2. From knowledge base to relational tables
@@
Information in data lakes is typically stored as tuples in tables, rather than the triplet format used in YAGO. For this reason, we construct YADL by first reshaping the triplets into binary tables. Wide-form relational tables can then be constructed by joining binary tables on the column that contains the triplet subjects. We assemble two types of YADL variants (Binary and Wordnet-based), which differ in the number of tables and in their shape (i.e., the number of rows and columns of each table).
@@
A binary table is generated for every predicate in YAGO, e.g., “hasCapital”. YAGO contains 70 predicates, which result is a small data lake where some of the tables have millions of rows (e.g., “isLocatedIn”) and some have few (e.g., “hasTLD”). Each table contains an attribute named “subject” and another attribute named as the actual predicate. For example, the triplet “France - hasCapital - Paris” leads to a table “hasCapital” with columns “subject” and “hasCapital”; finally, a row “France, Paris” is added to the table. The same process is applied to all 70 relations and their triplets.
Table 1. “City” seed table built by flattening predicates.
@@
Table 2. “City” seed table built by left-joining binary tables.
@@
To reflect that tables typically have more than two columns, we develop a second variant of YADL where tables have a larger number of columns. To this goal, we leverage the Wordnet (Miller, 1994) classes (e.g., “Person”, “Company”, “Artist”) to which entities belong. We use all Wordnet classes (a total of 1015) to create our seed tables.
For each class, a new table is generated following the example in Table 1. Initially, the seed table includes solely the “Subject” column (e.g., the “City” table includes only values “Paris,” “Rome,” etc.). Subsequently, we join every relation associated with subjects in the table: in Table 1, the subjects with type “City” are joined with relations “locatedIn”, “owns”, “hasPopDensity”.
@@
A subject may be linked to many objects through the same predicate, e.g., “Paris” is linked to two objects (“France”, “Europe”) via the predicate “locatedIn”. We address the issue in three ways. For Binary, triplets that share the same subject and predicate are transformed into tuples with the same value in the “subject” column and different values in the predicate column, e.g., the binary table has columns “Subject” and “isLocatedIn” and it contains tuples (Paris, France) and (Paris, Europe). For “Wordnet” seed tables, we build two variants. In the first case (which we dub YADL Base in section 5), we flatten the group of objects by creating new columns, thus moving them on the same tuple rather than splitting them (e.g., in Table 1, column “locatedIn” is flattened over “locatedIn_1” and “locatedIn_2”); in the second variant (YADL 10k and YADL 50k), we create the new columns by executing a left join between the subject and each binary table. One-to-many matches (such as with “Paris” and “France”, and “Paris” and “France”) are replicated with each left join; ordinarily, this would lead to an explosion in the number of rows: we limit the issue by selecting only the first two entries for each predicate (as shown in Table 2).
@@
For YADL Base, we generate all combinations of arity 2 and 3 for each seed table. For Table 1, these combinations would include “(isLocatedIn_1, owns_1)”, “(owns_1, popDensity)”, and “(isLocatedIn_1, owns_1, popDensity)”. Each sub-table is then built by projecting the seed table over the generated combination. We exclude all sub-tables with fewer than 100 rows. Rows containing null values are retained if the column combination includes at least one non-null value. For YADL 10k and 50k, we drop all seed tables whose arity is lower than a minimum arity $A$; then, for each surviving seed table $T$, we generate $N$ sub-tables by sampling a random subset of columns of size $[A-2,A]$ and projecting onto $T$; each parameter can be tweaked to modify the size and number of resulting sub-tables. Row-wise redundancy is provided by randomly sub-sampling a fraction $p$ of each resulting sub-table $n_{s}$ more times, thus replicating a fraction of the samples while keeping the set of columns fixed. We construct YADL 10k and YADL 50k by setting $N$ to 10 and 50 respectively, $A=8$, $p=0.7$, and $n_{s}=2$.
Data lake statistics are reported in Table 3, while Figure 2 reports the measured containment for the base tables we use in our experiments..  [^1] The data lakes are available at https://zenodo.org/doi/10.5281/zenodo.10600047, while the code to prepare YADL is at https://github.com/rcap107/YADL.
Figure 2. Exact containment measured over the top 100 join candidates for each base table (represented by a colored line) on each data lake. The x-axis reports the rank, rank 30 is the retrieval cutoff we use in our experiments.
4. Implementation of the pipeline
We now discuss the steps of the pipeline in Figure 1. For the Retrieve step, we describe the advantages and disadvantages of each candidate retrieval method; for the Merge step, we propose different join selectors and aggregation methods, and we explain how these two parts are intertwined; finally, for the Predict step, we go over the ML methods used to test the prediction performance.
4.1. Retrieving the candidates
We consider retrieval methods that work by taking a query column and return a – possibly ranked – list of candidates.
Retrieval methods usually involve offline preparation steps such as building sketches (Zhu et al., 2023; Fernandez et al., 2019), indices (Zhu et al., 2019), creating embeddings (Fan et al., 2023b), or measuring the Jaccard Containment of each column; preparation includes persisting the data structures for future use. Afterwards, candidate joins can be obtained by querying the retrieval method.
For retrieval strategies estimating Jaccard containment (Equation 1), a candidate column is a good join if it contains a large fraction of the entities of the query column (Section 5.2 provides an empirical justification). We consider two complementary strategies, an exact and a stochastic approximation, as well as a “hybrid” approach combining them. Starmie (Fan et al., 2023b) is our reference baseline.
The rest of the pipeline is transparent to the retrieval method: given a list of join candidates, it will test each candidate regardless of how the list was constructed. For our experiments, we reduce the number of candidates to the top-30, as we observe that almost all candidates with high containment are within this limit (Fig. 2).

### v6

@@ 3 Building Yet Another Data Lake
Yet Another Data Lake (YADL) is a semi-synthetic data lake built by recombining the data present in the YAGO (Suchanek et al., 2007; Mahdisoltani et al., 2014) knowledge base (KB)  [^1] We use YAGO 3.0.3 (Mahdisoltani et al., 2014), which is updated to 2022. to generate a collection of tables. Our goal is to have a scalable, high-quality and consistent data lake that allows users to evaluate the main steps of the pipeline in Figure 1, while avoiding some of the confounding factors that come from working with unvetted data, such as typos, inconsistent schemas and data format, and other sources of noise. These challenges can be added to YADL, e.g., by generating typos in table entries.
@@
YAGO: the source of the original data. YAGO (Suchanek et al., 2007) is a KB composed of RDF triplets; each triplet has a subject connected to an object through a predicate (or relation). Subjects and objects are considered to be entities, e.g., in the triplet “Paris, locatedIn, France”, “Paris” is the subject entity, “France” is the object entity, and “locatedIn” is the predicate that connects them. The object may also be a lexical value such as the one for the “population density”. Each entity belongs to a set of classes (or types), arranged in a taxonomy; “Paris” belongs to the class “City”, subclass of “Populated place”, itself subclass of “Geographical locations”.
From knowledge base to relational tables. Information in data lakes is typically stored as tables, rather than the triplet format of YAGO. For this reason, we first rearrange the YAGO triplets by converting them into binary tables, then we filter and recombine the binary tables to create multiple YADL variants (Base, 10k and 50k) that differ in their size and in the properties of the tables contained therein. YADL variants can be obtained by changing the parameters used for the synthesis, as detailed in Appendix A.
4 Implementations of the retrieve-merge-predict pipeline
We now discuss the pipeline steps in Figure 1. For the Retrieve step, we discuss different possible retrieval strategies; for the Merge step, we propose different join selectors and aggregation methods, and we explain how these two parts are intertwined; finally, for the Predict step, we go over the ML methods used to test the prediction performance. We explore each section in more detail in Appendix B.
Retrieving the candidates. We consider retrieval methods that work by taking a query column and return a – possibly ranked – list of candidates.
Retrieval methods usually involve offline preparation steps such as building sketches (Zhu et al., 2023; Fernandez et al., 2019), indices (Zhu et al., 2019), creating embeddings (Fan et al., 2023b), or profiling tables in the data lake (e.g., measuring the containment); preparation includes persisting the data structures for future use. Afterwards, candidate joins can be obtained by querying the retrieval method.
For retrieval strategies that rely on Jaccard containment (Equation 1), a candidate column is a good join if its containment is large, i.e. the candidate column contains a large fraction of the entities of the query column. Alternative strategies may rely on different metrics, such as cosine similarity between embeddings (Fan et al., 2023b).
We consider four retrieval methods that represent different levels of query cost to better observe the resource expense-effectiveness trade-off. Approximate methods are represented by MinHashLSHEnsemble (MinHash) (Zhu et al., 2016), which builds an index on the data lake, and when queried returns candidates whose Jaccard containment is greater than a certain threshold without ranking them. Conversely, Exact Matching measures the exact Jaccard containment between a query column and every column in the data lake. We propose Hybrid MinHash as a combination of Exact Matching and MinHash that takes the candidates returned by MinHash and measures their exact containment, thus limiting the number of columns to be considered and reducing the computational cost. Finally, Starmie (Fan et al., 2023b) builds embeddings for the query column and every column in the data lake using a language model, then combines cosine and Jaccard similarity to rank candidates. More detail is provided in Appendix subsection B.1.
The rest of the pipeline is transparent to the retrieval method: given a list of join candidates, it will test each candidate regardless of how the list was constructed. To reflect practical limitations on runtime and compute resources, we introduce a budget constraint on the number of retrieved candidates to evaluate. We choose to keep the top-30 candidates, as we observe that almost all candidates with high containment are within this limit (Figure 8 in Appendix), and it has empirical confirmation in practice (Spotify, 2020). In Section 5, we demonstrate how even such a small budget can cause problems with scalability.
Merging base table and join candidates. In this step, the retrieved candidates are combined into a new integrated table that joins the information in the base table $T$ with additional information from augmentation tables. As shown in Figure 1, the merge step combines join selection and aggregation, as both must be executed to build the augmented table used for Prediction.
Join selection
We consider two different overarching strategies for selecting candidates: metric-based selectors, which rely on a metric to choose which candidates should be joined, and results-based selectors, which instead iterate over candidates and select the best by explicitly executing joins. The first category includes Highest Containment, which ranks candidates by containment, then joins the first, and Top-$k$ Full Join, which relies entirely on the ranking provided by the retrieval method to join the first $k$ candidates without filtering them. For the second category, we implement Best Single Join and Stepwise Greedy Join. The first works with one candidate at each iteration, during which it first joins the candidate, then it trains and evaluates a prediction model on the resulting table; after iterating over all candidates, it selects the single candidate with the best performance. The second method iterates over each candidate like in the previous case, however it retains all candidates that improve the prediction performance, so that the augmented table grows over time as new candidates are joined. Stepwise Greedy Join is related to forward feature selection (Guyon & Elisseeff, 2003), as it greedily adds new tables as features during each iteration, and represents a common approach in the database literature (Galhotra et al., 2023; Chepurko et al., 2020). Appendix subsection B.2 details the implementations.
Aggregation
When joining tables for a downstream ML task, one-to-many relationships must be aggregated to avoid replicating samples in the base table, which would modify the initial data sampling. We test three aggregation strategies: Any selects any row at random from each group of matched tuples; Mean replaces all duplicated numerical (categorical) values by the mean (most frequent) of all values for that attribute in the group; finally DFS (Deep Feature Synthesis (Kanter & Veeramachaneni, 2015)) greedily generates new features for each column in the augmented table by aggregating groups of tuples, measuring statistics over the groups (e.g., mean, median, most frequent value), and adding said statistics as features; this is done for every join. We expand on this in Appendix subsection B.3.
Supervised learning with a ML model. The learning step is performed on the final integrated table on the training split returned by the join selector and evaluated on the integrated table on the test split.
We evaluate four predictors. We use a linear baseline: Ridge regressor/classifier with the RidgeCV scikit-learn (Pedregosa et al., 2011) implementation and default parameters. CatBoost (Prokhorenkova et al., 2018) is a state-of-the-art GBDT method, optimized to handle categorical variables. We set the number of iterations to 300. ResNet (Gorishniy et al., 2021) is our baseline neural method. RealMLP (Holzmüller et al., 2024) is a NN-based method that incorporates a number of “good defaults” to improve performance over standard NNs. For ResNet and RealMLP we use the implementation and default parameters provided by pytabkit (Holzmüller et al., 2024). We do not perform hyperparameter optimization.
5 Experimental study
5.1 Settings
For our experimental campaign, we test the different sections of the pipeline over three dimensions: prediction performance, execution time, and RAM usage. Focusing on multiple dimensions allows us to have a view of the different trade-offs across methods.
Data lakes
We use four YADL variants (Binary, Base, 10k, 50k) (as detailed in Section 3) and Open Data US, a data lake employed in the literature (Galhotra et al., 2023; Zhu et al., 2019; 2016); statistics are reported in Table 1. Starmie did not run on YADL 50k (not enough RAM) and Open Data (noise in the data crashed the model), so these data lakes are excluded from figures that involve Starmie. Appendix Figure 8 reports the measured containment for the base tables used in our experiments. [^2] The data lakes are available at https://doi.org/10.5281/zenodo.10600047, while the code to prepare YADL is at https://github.com/rcap107/YADL and the pipeline is at https://github.com/rcap107/retrieve-merge-predict.
Table 1: Statistics for the tables in the data lakes. “N. cols” and “C. cols” refer to numerical and categorical columns, respectively.
YADL YADL OpenData
Binary Base 10k 50k US
N. tables 70 30072 10059 47223 5591
Size (MB) 629 10051 5557 27718 4062
Tot. rows 20.1M 672M 242M 1.20B 95.7M
Tot. cols 140 95.2k 127k 624k 133k
Avg rows 287k 22.3k 24.0k 25.4k 17.1k
Avg cols 2.00 3.17 12.63 13.21 23.86
Avg N. cols 0.30 0.39 3.54 3.59 11.10
Avg C. cols 1.70 2.78 9.08 9.61 12.76
Avg nulls 0.00 0.31 0.62 0.64 0.09
Base tables
We evaluate five tables from sources external to the lakes: Company Employees, US Elections, 2021 US Road Accidents, and Housing Prices. As “internal” tables, derived from lakes, have been used in several previous works for evaluation (Galhotra et al., 2023; Chepurko et al., 2020), we also include one per lake: US County Population (from YADL), and Schools (from Open Data US). Statistics are reported in Appendix subsection C.8 .
Figure 2: Pareto frontier analysis for the pipeline steps The prediction performance is plotted against run time (top) and peak RAM usage (bottom). Each row presents the same results, broken down by retrieval method (left), join selector method (center) and predictor (right). Each dot represents the average prediction performance and resource cost averaged across base tables and data lakes for a specific configuration (e.g., the leftmost dot in the first row is obtaining by using MinHash, Highest Containment Join and CatBoost). Time for offline retrieval preparation, i.e. indexing and querying, is not reported here.
For all datasets, the values of the query columns must be matched with the entities in YADL using semantic annotation solutions (Huynh et al., 2022; Nguyen et al., 2021). In our experiments, we manually performed the match to remove the noise from this task. Also, query columns are chosen based on what a user may reasonably consider as “key” (e.g., the movie title). We release the matched tables along with the data lakes for reproduciblity.
To reflect the experience of a data scientist that needs to construct a meaningful table starting from a data lake, and to highlight the effect of joins on the downstream task, we run experiments on a depleted version of the tables, i.e., the input tables include only the primary key column and the target column. We provide details on the datasets, computational resources and pipeline implementation in Appendix subsection C.2.
5.2 Results
Our first goal is to pinpoint which steps of the pipeline have the most significant impact on the studied dimensions: optimizing these influential sections can yield the greater benefits. We rely on Pareto frontier plots (Figure 2) to highlight the cost-performance trade-off for the different steps and configurations.
Reference configuration
Based on Figure 2, we define a reference configuration that uses Exact Matching, Best Single Join, aggregation Any, and CatBoost. This configuration represents a good trade-off between performance and compute cost, and it can run on all configurations, base tables and data lakes.
Predictor: Tree-based methods perform well
Supervised learning with the ML model is the pipeline step with the starkest difference between methods (Figure 2 right, Table 3): in our scenario, CatBoost is both faster and more effective than its non-tree counterparts. Indeed, CatBoost always outperforms other methods in prediction performance (by up to 26%) and run-time ($\times3$ faster than the slowest competitor), although it has a larger memory footprint.
This is likely due to two major factors: 1) imperfect joins introduce a large fraction of missing values (database “nulls”), and 2) categorical features have high cardinality. Firstly, any sample that does not find a match in a candidate will have a missing value in each new feature; in addition, any missing value in matched rows will remain: as a result, even “good” joins may feature a high degree of missingness. Parametric models are particularly affected by missing values, and categorical features make imputation more difficult; comparatively, tree-based models are far more resilient in presence of missing values (Josse et al., 2024). In fact, we observe that CatBoost and RealMLP are the only methods whose prediction performance never drops below 0 even in this very challenging scenario. McElfresh et al. (2023) also report that heavy missingness impedes neural networks more than tree-based models.
Figure 3: Retrieval: Better containment improves prediction performance. Regression plot relating the prediction performance with the Jaccard containment; each dot represents an experimental run for each base table.
Retrieval: Containment is key, though imperfect
A surprising result of our experiments shown in Figure 2 and Table 3 is that simple metric-based retrieval is sufficient to achieve good results, while maintaining a far lower resource cost. Relying directly (or indirectly) on Jaccard containment (Highest Containment, Full Join, Exact Matching) achieves the same results as far more thorough (and costly) retrieval (Starmie) or selection (Best Single Join, Stepwise Greedy Join).
Focusing on the reference configuration, Figure 3 relates containment to prediction performance. The plotted lines highlight the clear upwards trend in the prediction performance that follows the increase in containment. In other words, higher prediction results occur more frequently when the containment is higher. For exact joins (as in our scenario), having a high Jaccard containment leads to better results because left joins with high overlap augment a large fraction of the base table, whereas low overlap joins produce features that contain mostly empty values.
And yet, while Jaccard containment is an effective first metric, it is not fool proof. We observe experimentally that it does not work well where there is a large degree of redundancy because in such cases many candidates may share the same containment: this is what happens with YADL 50k (Table 2, Appendix Figure 8); a larger computational budget may mitigate this issue, but it is not always an option. Furthermore, it fails in cases where the cardinality of a column is very low (e.g., with columns that contain binary values): this is problematic if the user does not know which columns should be used for querying. In other words, high containment means that we can join without adding too many missing values: this does not guarantee that the added features improve downstream utility.
Retrieval: Better metrics improve performance, at a cost
Pareto frontier analysis (Figure 2, left) shows that Starmie is Pareto optimal for some (resource-intensive) configurations: if resources are available for the offline preparation cost, better querying through Starmie can bring benefits. And yet, the –arguably small– improvements in downstream performance brought by Starmie come with a very large cost in both RAM and offline training time (Figure 5, Appendix Table 6, Appendix Table 10). Furthermore, we had to restrict most of our experiments (including Figure 2) as Starmie could only run on 3 out of our 5 data lakes of interest (Binary, YADL Base and YADL 10k). We had to leave out YADL 50k, where Starmie exceeded our available RAM (150GB), and Open Data which crashed Starmie, probably because it was too noisy. On the other hand, Exact matching is Pareto optimal in most cases, and runs on all data lakes we considered.
With respect to MinHash, Hybrid MinHash brings a substantial improvement in terms of downstream performance (Figure 2, Table 3, Figure 15 ) and top-1 retrieval performance (Appendix Figure 18, Appendix Figure 16). It leverages Exact Matching in order to produce an actual ranking of the candidates. This however comes at the cost of a substantial increase in the query time (Figure 5, Appendix Table 10).
Figure 4: Resource cost of indexing the retrieval methods. The computational cost of each retrieval method is plotted with respect to the reference (Exact Matching). The median difference is reported on the right of each plot.
Figure 5: Runtime as a function of the number of query columns for the different retrieval methods.
Retrieval: Computation trade-offs
Figure 5 compares resources for the different retrieval methods at indexing and query time [^3] We separate retrieval proper from the pipeline execution to clearly highlight their corresponding cost and trade-offs.: Starmie is by far the most expensive method, both in run time and in peak RAM. The rapid growth of RAM usage, even for relatively small data lakes (see Appendix Table 6), is particularly problematic and prevents the method from running on GPUs, thus slowing it down even further. The other methods are far cheaper by comparison, and can run on CPU without issues.
Figure 5 shows how different retrieval methods scale with the number of query columns: the value for 0 columns shows the time required to build the index (averaged over all data lakes [^4] Open Data and YADL 50k not included.), and the slope is the time required to query a single column (averaged over all query columns); values are reported in Appendix Table 10. The figure highlights the point at which the cost of recomputing the containment for each new query column on every column in the data lake (Exact Matching) becomes more expensive than building the index (MinHash), and then querying it.
Figure 6: Aggregation: Pareto Frontier Analysis. DFS outperforms other methods, but is much slower to run.
It is clear that, as the number of query columns increases, Exact Matching scales much worse than the other solutions (except Starmie), while Hybrid MinHash and MinHash scale much better: this is important if the query column is unknown or a user would like to test multiple columns to augment one or more tables. In any case, Starmie is much slower than all other alternatives.
Aggregation: Complex aggregation improves performance with limitations
Figure 6 reports aggregation results [^5] More detail is available in Appendix, Figure 13 and Figure 14. The systematic expansion of multiple aggregations in DFS leads to many features, which would make both Full Join and Stepwise Greedy Join intractable, so they are not considered here. The generated features overall bring useful information, as DFS outperforms the simpler aggregations in most cases; however, preparing and employing the new features increases noticeably the training time. This is confirmed by Table 3, which highlights both the improvement in performance ( 2.5%) and the major increase of compute cost ( 6x wrt Any).
Overall results and ablation study
Table 2: Aggregated median prediction results (higher is better) for the reference configuration. “NA” results are referring to experiments run on “internal” tables (“Schools” was sampled from Open Data, “US County Population” from YADL): if run on a different data lake, no matching entities could be found, so no new features would be added.
Base Table Binary YADL Base YADL 10k YADL 50k Open Data US
Company Employees 0.20 0.33 0.37 0.25 0.26
Housing Prices 0.34 0.57 0.54 0.54 0.50
Schools NA NA NA NA 1.00
2021 US Accidents 0.26 0.44 0.44 0.43 0.31
US County Population 0.93 0.84 0.95 0.85 NA
US Elections 0.44 0.55 0.52 0.52 0.59
Table 3: Ablation study, prediction difference (higher is better) and runtime difference (lower is better) from the reference configuration. Retrieval Method does not include results from Open Data US and YADL 50k.
Retrieval Method Prediction diff. Runtime diff. Join Selector Prediction diff. Runtime diff.
Exact Matching 0.00% 1.00x Full Join 2.09% 0.57x
Starmie -0.06% 0.98x Stepwise Greedy Join 1.61% 2.91x
Hybrid Minhash -3.19% 0.67x Best Single Join 0.00% 1.00x
Minhash -17.48% 0.44x Highest Containment Join -0.10% 0.12x
Aggregation Prediction diff. Runtime diff. Predictor Prediction diff. Runtime diff.
DFS 2.46% 6.08x Catboost 0.00% 1.00x
Mean 0.02% 1.03x RealMLP -21.24% 32.79x
Any 0.00% 1.00x ResNet -24.11% 10.28x
RidgeCV -26.03% 5.98x
Table 2 reports the median prediction performance across all configurations for a pair “Base table - Data Lake”. From this table it is evident that internal tables (Schools, US County Population) represent a much simpler problem to solve compared to tables that are not sampled from the data lake (all other tables). Table 3 reports the results of an ablation study where we compare the performance and runtime of the reference configuration against the other possible methods; we report the median difference in prediction and runtime.  [^6] This is expanded in Appendix subsection C.2 and Figure 15.
6 Discussion and conclusions
Take-away messages
We build a semi-synthetic data lake based on a knowledge base to use as a reproducible test bed for evaluating methods to augment user-provided tables. We implement an easily-extendable pipeline to test the different steps of the augmentation procedure and alternative algorithms for each of their tasks. Our results uncover a number of observations important to direct research on this subject. We summarize the main take-away messages from the experiments.
  1. 1.
    Tree-based models come with significant benefits in terms of prediction and computational performance in the data-lake pipeline we studied (Figure 2 right, Table 3). Indeed, the automated selection of augmentation candidate tables generates challenging features (e.g., with many missing values), which tree-based models are more effective at dealing with.
  2. 2.
    Good table retrieval affects the entire pipeline by discovering candidates that contain more useful features and introduce fewer missing values (Figure 2 left). Jaccard containment is a good metric for retrieving candidates, but it has limitations (Figure 3).
  3. 3.
    Simple metric-based retrieval (Exact matching) and candidate selection (Highest containment) produce comparable or better results than more complex methods (Starmie [^7] Note that our results are not in contradiction with the Starmie study, as it was focused on table unionability, i.e. finding more samples rather feature augmentation which is our focus. and Best single join), while being vastly more efficient (Figure 2 left, center, Table 3).
  4. 4.
    Complex aggregation methods are much slower than simpler ones and do not result in commensurate gains in prediction performance (Figure 6, Table 3).
  5. 5.
    Combining stochastic retrieval with exact metrics (Hybrid MinHash) mitigates some of the drawbacks of basic MinHash and scales well with many query columns (Figure 3(a), Figure 5, Figure 5, Table 3).
The topic that we have studied is at the intersection of database and machine learning research. This intersection still has much open research prospect:
Merging on clean columns is only one side of the story. While we focus on columns where join keys can be matched exactly (functionally, by doing string matching), this is not possible in general because of typos, different formats, and different granularity. Similarity joins and semantic matching would help with this problem (Jiang et al., 2014; Dong et al., 2023; Deng et al., 2017b; Mundra et al., 2023) and add another dimension to the analysis. Methods that cover multiple steps, such as (Gong et al., 2023; Huang et al., 2023c), could also be considered.
To limit the search space, we keep our chain of joins for augmentation limited to one, and show that scalability is an issue even in this simplified scenario. However, some methods have considered chains of join for augmentation (Galhotra et al., 2023). Enabling join chains would also make evident the benefit of recursive methods (Cvetkov-Iliev et al., 2023; Kanter & Veeramachaneni, 2015).
“AutoRetrieval” is an exciting prospect. The overall middling-to-poor prediction results (Table 2, Table 8) suggest that the automated strategies evaluated here are not sufficient to replicate the performance achieved by a human data analyst. Automated procedures may help with discovering good candidates, but they cannot replace a human expert: fully automating this operation similarly to how AutoML tools (Hutter et al., 2019; Feurer et al., 2015) tweak ML methods represents a compelling direction for future work.
Final words. The performance-complexity trade-offs that appear are crucial for the practitioner: they suggest how to automate supervised learning on a data lake of a given size, maximizing statistical performance within a compute budget.
Appendix A Construction of YADL
@@
We construct YADL by first reshaping the RDF triplets into binary tables. Then, we group subjects by their Wordnet type and construct 1015 wide-form relational “seed tables” by joining binary tables on the column that contains the triplet subjects.
Each seed table is then partitioned into “sub-tables” by selecting subsets of columns and projecting over them. Different partitioning strategies allow us to vary the characteristics of the resulting sub-tables (i.e., the number of rows and columns of each table), and the size of the resulting data lake.
@@
A binary table is generated for every predicate in YAGO, e.g., “hasCapital”. YAGO contains 70 predicates, which result in a small data lake where some of the tables have millions of rows (e.g., “isLocatedIn”) and some have few (e.g., “hasTLD”). Each table contains an attribute named “subject” and another attribute named as the actual predicate. For example, the triplet “France - hasCapital - Paris” leads to a table “hasCapital” with columns “subject” and “hasCapital”; finally, a row “France, Paris” is added to the table. The same process is applied to all 70 relations and their triplets.
Table 4: “City” seed table built by flattening predicates.
@@
Table 5: “City” seed table built by left-joining binary tables.
@@
To reflect that tables typically have more than two columns, we develop a second variant of YADL where tables have a larger number of columns. To this goal, we leverage the Wordnet (Miller, 1994) classes (e.g., “Person”, “Company”, “Artist”) to which entities belong. We use all Wordnet classes (a total of 1015) to create our seed tables, following the example in Table 4. Initially, the seed table includes solely the “Subject” column (e.g., the “City” table includes only values “Paris,” “Rome,” etc.). Then, we join every relation associated with subjects in the table: in  Table 4, the subjects with type “City” are joined with relations “locatedIn”, “owns”, “hasPopDensity”.
@@
A subject may be linked to many objects through the same predicate, e.g., “Paris” is linked to two objects (“France”, “Europe”) via the predicate “locatedIn”. We address the issue in three ways. For Binary, triplets that share the same subject and predicate are transformed into tuples with the same value in the “subject” column and different values in the predicate column, e.g., the binary table has columns “Subject” and “isLocatedIn” and it contains tuples (Paris, France) and (Paris, Europe). For “Wordnet” seed tables, we build two variants. In the first case (which we dub YADL Base in Section 5), we flatten the group of objects by creating new columns, thus moving them on the same tuple rather than splitting them (e.g., in Table 4, column “locatedIn” is flattened over “locatedIn_1” and “locatedIn_2”). In the second variant (YADL 10k and YADL 50k), we create the new columns by executing a left join between the subject and each binary table. One-to-many matches (such as with “Paris” and “France”, and “Paris” and “Europe”) are replicated with each left join; ordinarily, this would lead to an explosion in the number of rows: we limit the issue by selecting only the first two entries for each predicate (as shown in Table 5).
@@
For YADL Base, we generate all combinations of arity 2 and 3 for each seed table. For Table 4, these combinations would include “(isLocatedIn_1, owns_1)”, “(owns_1, popDensity)”, and “(isLocatedIn_1, owns_1, popDensity)”. Each sub-table is then built by projecting the seed table over the generated combination; we retain rows that contain at least one non-null value and drop sub-tables with fewer than 100 rows. For YADL 10k and 50k, we drop all seed tables whose arity is lower than a minimum arity $A$; then, for each surviving seed table $T$, we generate $N$ sub-tables by sampling a random subset of columns of size $[A-2,A]$ and projecting onto $T$; each parameter can be tweaked to modify the size and number of resulting sub-tables. Row-wise redundancy is provided by randomly sub-sampling a fraction $p$ of each resulting sub-table $n_{s}$ more times, thus replicating a fraction of the samples while keeping the set of columns fixed. We construct YADL 10k and YADL 50k by setting $N$ to 10 and 50 respectively, $A=8$, $p=0.7$, and $n_{s}=2$.
Appendix B Detailed description of the tested methods
B.1 Retrieval